OpenStudy (anonymous):

Tomas bought 3 software apps and 2 e-books for a total of $21. The two e-books cost $12. If s represents the cost of each app, the total cost of all 5 items is represented by 3s+12=21 . Which statement about verifying the solution for s is true?

OpenStudy (anonymous):

@chunkymonkey

OpenStudy (anonymous):

@julieishere21

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

I need to see the statements

OpenStudy (anonymous):

3s + 12 = 21 (subtract 12 from both sides) 3s = 21 - 12 3s = 9 (divide both sides by 3) s = 3

OpenStudy (anonymous):

A regular hexagon and a rectangle have the same perimeter, P. A side of the hexagon is 4 less than the length, l, of the rectangle. The width of the rectangle, w, is 2 less than the length of the rectangle. What is the perimeter of the hexagon?

OpenStudy (anonymous):

I am sorry.....I do not know this one

OpenStudy (anonymous):

Kendall scored 3 times as many points as Abdul scored in a basketball game. Otis scored 5 more points than Abdul scored. The three players scored 45 points all together. How many points did each player score? Kendall: 24 points; Abdul: 13 points; Otis: 8 points Kendall: 8 points; Abdul: 8 points; Otis: 8 points Kendall: 24 points; Abdul: 8 points; Otis: 13 points Kendall: 8 points; Abdul: 13 points; Otis: 24 points

OpenStudy (anonymous):

K = 3A O = A + 5 K + O + A = 45 (3A) +( A + 5) + A = 45 5A + 5 = 45 5A = 40 A = 8 K = 3A K = 3(8) = 24 O = A + 5 O = 8 + 5 O = 13 check... K + O + A = 45 24 + 13 + 8 = 45 45 = 45 (correct) Kendall = 24, Abdul = 8, Otis = 13

OpenStudy (anonymous):

Amie works 6 hours each Saturday and 4 hours each Monday. She also makes $20 a week as a dog walker. She makes a total of $130 each week. If h represents Amie’s hourly rate, which equation can be used to find Amie’s hourly rate? 6h+20=130 (6+4)h=30 (6+4)h+20=130 4h+20=130

OpenStudy (anonymous):

btw your really smart&Thanks!

OpenStudy (anonymous):

(6 + 4)h + 20 = 130

OpenStudy (anonymous):

wait...let me look at this again

OpenStudy (anonymous):

ok i dont think i need help with anything else i should get a 70%

OpenStudy (anonymous):

the dog walking should not effect her hourly rate.... but none of the other answers makes sense...

OpenStudy (anonymous):

I am thinking it could be (6 + 4)h = 30

OpenStudy (anonymous):

not 100% sure on this one

OpenStudy (anonymous):

In a basketball tournament, team A scored 6 more points than 3 times as many points as team B scored. Team C scored 45 more points than team B scored. The combined scores for all three teams was 476 points. Which statement about the number of points is true? Team A scored 85 points. Team A scored 255 points. Team C scored 261 points. Team C scored 130 points.

OpenStudy (anonymous):

last one(: